40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14022     Entry time: Tue Jun 26 20:59:36 2018
Author: aaron 
Type: Update 
Category: OMC 
Subject: prep for vent in a couple weeks 

I checked out the elog from the vent in October 2016 when the OMC was removed from the path. In the vent in a couple weeks, we'd like to get the beam going through the OMC again. I wasn't really there for this last vent and don't have a great sense for how things go at the 40m, but this is how I think the procedure for this work should approximately go. The main points are that we'll need to slightly translate and rotate OM5, rotate OM6, replace one mirror that was removed last time, and add some beam dumps. Please let me know what I've got wrong or am missing.

[side note, I want to make some markup on the optics layouts that I see as pdfs elsewhere in the log and wiki, but haven't done it and didn't much want to dig around random drawing software, if there's a canonical way this is done please let me know.]

Steps to return the OMC to the IFO output:

  1. Complete non-Steve portions of the pre-vent checklist (https://wiki-40m.ligo.caltech.edu/vent/checklist)
  2. Steve needs to complete his portions of the checklist (as in https://nodus.ligo.caltech.edu:8081/40m/12557)
  3. Need to lock some things before making changes I think—but I’m not really sure about these, just going from what I can glean from the elogs around the last vent
    1. Lock the IMC at low power
    2. Align the arms to green
    3. Lock the arms
    4. Center op lev spots on QPDs
    5. Is there a separate checklist for these things? Seems this locking process happens every time there is a realignment or we start any work, which makes sense, so I expect it is standardized.
  4. Turn/add optics in the reverse order that Gautam did
    1. Check table leveling first?
    2. Rotate OM5 to send the beam to the partially transmissive mirror that goes to the OMC; currently OM5 is sent directly to OM6. OM5 also likely needs to be translated forward slightly; Gautam tried to maintain 45 deg AOI on OM5/6.
    3. A razor beam dump was also removed, which should be replaced (see attachment 1 on https://nodus.ligo.caltech.edu:8081/40m/12568)
    4. May need to rotate OM6 to extract AS beam again, since it was rotated last time
    5. Replace the mirror just prior to the window on the AP table, mentioned here in attachment 3: https://nodus.ligo.caltech.edu:8081/40m/12566
      1. There is currently a rectangular weight on the table where the mirror was, for leveling
  5. Since Gautam had initially made this change to avoid some backscattered beams and get a little extra power, we may need to add some beam dumps to kill ghosts
    1. This is also mentioned in 12566 linked above, the dumps are for back-reflection off the windows of the OMC
  6. Center beam in new path
  7. Check OMC table leveling
  8. AS beam should be round on the camera, with no evidence of clipping on any optics in the path (especially check downstream of any changes)
ELOG V3.1.3-